# SatFed: A Resource-Efficient LEO Satellite-Assisted Heterogeneous Federated Learning Framework

### v2

@@ I Introduction
As edge devices like drones, smart cameras, and IoT sensors proliferate, they generate vast amounts of data at the network edge [1]. This data drives significant advancements in key domains such as video analytics [2], resource management [3], and autonomous driving [4] through machine learning (ML). However, privacy considerations [5] and bandwidth constraints [6] render transferring large volumes of data to a cloud server impractical, necessitating the adoption of distributed ML training methods.
To address these challenges, federated learning (FL) [7] has emerged as a promising distributed machine learning framework, allowing edge devices to collaboratively train models without sharing local data. FL typically employs a parameter server (PS) architecture, where a central server aggregates model updates from participating devices. The efficiency of FL training is heavily dependent on the quality of communication, given the continuous exchange of large volumes of parameters. However, existing FL systems solely rely on terrestrial networks, which are increasingly strained by rising user numbers and service demands [8, 9]. This often results in delays or congestion in parameter transmission [10], significantly affecting FL convergence, especially in remote or underdeveloped regions with poor signal coverage [11].
@@
Fortunately, the proliferation of low Earth orbit (LEO) satellites is revolutionizing satellite broadband Internet [12, 13], providing a crucial solution to the limitations of traditional terrestrial FL. Specifically, SpaceX plans to deploy about $42,000$ LEO satellites for its Starlink constellation [14]. LEO satellite networks offer broad coverage, facilitating peer-to-peer communication among numerous edge devices [15]. For example, when devices engaged in FL experience unstable terrestrial connections with central servers, satellite connectivity allows them to rapidly obtain knowledge from multiple peers, thereby improving their local training.
@@ II-B Data Heterogeneity
The performance of satellite-assisted FL is significantly impacted by data heterogeneity, as devices distributed globally encounter data from diverse contexts. These devices exhibit non-independent and non-identically distributed (non-IID) data, resulting in varied local optimization objectives [21]. Consequently, model updates from different devices are biased toward their local data distributions. Directly averaging these updates, as is done in the widely used FL method FedAvg [7], often leads to suboptimal performance. To better understand the impact of data heterogeneity on training, we conducted experiments with the ResNet-50 model using FedAvg at varying levels of heterogeneity, as shown in Fig. 3(a). The results clearly indicate that as data heterogeneity increases, performance degrades significantly.
@@ VI Conclusion
In this paper, we propose SatFed, a resource-efficient satellite-assisted FL framework designed to enhance training performance and convergence speed in complex heterogeneous environments. Initially, SatFed designs a priority transmission queue based on model freshness differences, enabling devices in satellite networks with highly constrained transmission rates and complex topologies to receive near real-time peer models. Furthermore, SatFed utilizes a multigraph to perceive in real-time the heterogeneity relationships among devices implied by models transmitted in satellite networks. Finally, SatFed utilizes perceived heterogeneity relationships to help devices in constructing a peer guide model from satellite transmission, directing local updates. Extensive experiments show that the SatFed framework outperforms state-of-the-art benchmarks. This work underscores the potential of deploying SatFed across real LEO satellite-assisted terrestrial networks.

### v3

@@ I Introduction
As edge devices like drones, smart cameras, and IoT sensors proliferate [1, 2, 3], they generate vast amounts of data at the network edge [4, 5]. This data drives significant advancements in key domains such as resource management [6, 7], authentication[8], image compression [9], and autonomous driving [10, 11, 12, 13] through machine learning (ML). However, privacy considerations [14, 15] and bandwidth constraints [16] render transferring large volumes of data to a cloud server impractical, necessitating the adoption of distributed ML training methods.
To address these challenges, federated learning (FL) [17] has emerged as a promising distributed machine learning framework, allowing edge devices to collaboratively train models without sharing local data [18, 19, 20]. FL typically employs a parameter server (PS) architecture, where a central server aggregates model updates from participating devices. The efficiency of FL training is heavily dependent on the quality of communication, given the continuous exchange of large volumes of parameters [21]. However, existing FL systems solely rely on terrestrial networks, which are increasingly strained by rising user numbers and service demands [22, 23, 24]. This often results in delays or congestion in parameter transmission [25], significantly affecting FL convergence, especially in remote or underdeveloped regions with poor signal coverage [26].
@@
Fortunately, the proliferation of low Earth orbit (LEO) satellites is revolutionizing satellite broadband Internet [27, 28, 29, 30, 31], providing a crucial solution to the limitations of traditional terrestrial FL. Specifically, SpaceX plans to deploy about $42,000$ LEO satellites for its Starlink constellation [32]. LEO satellite networks offer broad coverage, facilitating peer-to-peer communication among numerous edge devices [33]. For example, when devices engaged in FL experience unstable terrestrial connections with central servers, satellite connectivity allows them to rapidly obtain knowledge from multiple peers, thereby improving their local training.
@@ II-B Data Heterogeneity
The performance of satellite-assisted FL is significantly impacted by data heterogeneity, as devices distributed globally encounter data from diverse contexts [39]. These devices exhibit non-independent and non-identically distributed (non-IID) data, resulting in varied local optimization objectives [40]. Consequently, model updates from different devices are biased toward their local data distributions. Directly averaging these updates, as is done in the widely used FL method FedAvg [17], often leads to suboptimal performance. To better understand the impact of data heterogeneity on training, we conducted experiments with the ResNet-50 model using FedAvg at varying levels of heterogeneity, as shown in Fig. 3(a). The results clearly indicate that as data heterogeneity increases, performance degrades significantly.
@@ VI Conclusion
In this paper, we propose SatFed, a resource-efficient satellite-assisted FL framework designed to enhance training performance and convergence speed in complex heterogeneous environments. Initially, SatFed designs a priority transmission queue based on model freshness differences, enabling devices in satellite networks with highly constrained transmission rates and complex topologies to receive near real-time peer models. Furthermore, SatFed utilizes a multigraph to perceive in real-time the heterogeneity relationships among devices implied by models transmitted in satellite networks. Finally, SatFed utilizes perceived heterogeneity relationships to help devices in constructing a peer guide model from satellite transmission, directing local updates. Extensive experiments show that the SatFed framework outperforms state-of-the-art benchmarks. In the future, we can explore more advanced distributed learning paradigms such as split learning [57, 58], and more powerful large language model [59, 60, 61, 62] deployments in satellite networks.